# Spectral results for free random variables

### v1

@@ 5.1. The case of $ub_{t}$
According to [8, Theorem 2.7], $S$ satisfies the PDE
@@
For all $t>0,$ the spectrum of $ub_{t}$ is contained in $\overline{\Sigma}_{t}.$ Thus, since [23, Theorem 4.28] shows that $\mathrm{supp}(\mathrm{Br}_{ub_{t}})=\overline{\Sigma}_{t}$ and since the support of the Brown measure of any element is contained in its spectrum, we conclude that
@@
If we evaluate at $\varepsilon_{0}=0,$ the lifetime of the solution of the Hamiltonian system is $T(\lambda),$ which is greater than $t$ by Lemma 5.3. Then a general result about flows (e.g., the fact that the sets $M_{t}$ in [25, Theorem 9.12] are open) tells us that for $(\lambda_{0},\varepsilon_{0})$ in a neighborhood of $(\lambda,0),$ the lifetime of the solution with initial conditions $(\lambda_{0},\varepsilon_{0})$ will remain greater than t.t. [^1] This point is more subtle than it may appear because the formula for the lifetime of the Hamiltonian system in [8, Proposition 5.9] is only valid for $\varepsilon_{0}\geq 0.$ The difficulty is that when $\varepsilon_{0}<0,$ the blowup time of the whole system may be smaller than the blowup time in the formula for $p_{\varepsilon}(t)$. See Remark 5.10 in [8].
@@
from [8, Equation (5.8)], where the notation means that $p_{\varepsilon}$ is computed using the initial conditions $\lambda_{0}$ and $\varepsilon_{0}.$ Then
@@
As we have noted, $T(\lambda)$ has a removable singularity at $\left|\lambda\right|=1.$ Furthermore, $p_{\varepsilon,0}(\lambda)$ can be infinity but cannot be zero. Thus, the only way $T(\lambda)$ can be infinite is when $\lambda=0.$ Meanwhile, according to Proposition 4.8 in [23], the function $T(\lambda)$ is the decreasing limit as $\varepsilon_{0}\rightarrow 0^{+}$ of a certain function $t_{\ast}(\lambda,\varepsilon_{0})$ and this function is continuous in $\lambda$ for each $\varepsilon_{0}.$ The claimed semicontinuity of $T$ then follows by an elementary result [30, Theorem 15.84] about semicontinuous functions. ∎
@@
The claimed result is stated in Theorem 4.10 of [23]. There is, however, a small gap in the proof, concerning the case $\left|\lambda\right|=1,$ which we fill in here. Since $T(0)=\infty,$ we only consider nonzero $\lambda.$ Now, a point $\lambda$ with $T(\lambda)=t>0$ will be outside $\overline{\Sigma}_{t}$ if and only if $T$ has a weak local minimum at $\lambda$ (meaning that all $\lambda^{\prime}$ in a neighborhood of $\lambda$ have $T(\lambda^{\prime})\geq t,$ so that such points are outside $\Sigma_{t}$). According to [23, Lemma 4.15], the function $T(re^{i\theta}),$ with $\theta$ fixed, is strictly increasing for $1<r<\infty,$ and strictly decreasing for $0<r<1.$ Thus, any possible weak local minimum of $T$ would have to be at a point on the unit circle.
Consider, then, a point $e^{i\theta}$ on the unit circle that is outside $\overline{\Sigma}_{t}$ and thus (Lemma 5.2) outside $\sigma(u)=\mathrm{supp}(\mu_{u}).$ Putting $\left|\lambda\right|=1$ in (5.4), we may then compute that
@@
Thus, $1/T$ cannot have a weak local maximum at $e^{i\theta}$ and $T$ cannot have a weak local minimum at $e^{i\theta}.$ (Note, however, that $T$ can have a weak local minimum on the unit circle, namely when it is zero—in which case, (5.6) becomes meaningless—but this cannot happen at points in the unit circle outside $\overline{\Sigma}_{t}.$) ∎
@@ 5.2. The case of $ub_{t,\gamma}$
where $b_{t}$ is the value of $b_{t,\gamma}$ at $\gamma=0.$ By the $\tau=s$ case of [18], this map agrees with the one denoted $\Phi_{s,\tau}$ in [18, Section 8]. The following result shows that the “model deformation phenomenon” holds in this setting. That is to say, as we vary $\gamma$ with with $u$ and $t$ fixed, the Brown measure of $ub_{t,\gamma}$ changes in a very specific way, namely by push-forward under the map $\Psi_{t,\gamma}.$
@@
This result is due to Ho–Zhong [23, Corollary 4.30] in the case $\gamma=0$ and to Hall–Ho [18, Theorem 8.2] for $\gamma\neq 0.$
@@
Thus, $\Psi_{t,\gamma}$ coincides on $(\overline{\Sigma}_{t})^{c}$ with the holomorphic function denoted $f_{\gamma}$ in [18, Definition 2.2]. Furthermore, $\Psi_{t,\gamma}$ is defined and continuous on $\mathbb{C}.$
@@
The formula (5.10) follows from the $\tau=s$ case of [18, Theorem 6.1]. Once (5.10) is established, we see that $\Psi_{t,\gamma}$ coincides on $(\overline{\Sigma}_{t})^{c}$ with the function denoted $f_{s-\tau}$ in [18], where $s$ and $\tau$ in [18] correspond to $t$ and $t-\gamma$, respectively, here. Then the claimed injectivity follows from Theorem 3.8 in [18]. Note that this theorem assumes $\tau\neq 0,$ which means $\gamma\neq t$ in our current notation, but in light of Lemma 5.3, this assumption is only needed to ensure injectivity of $f_{s-\tau}$ on the boundary of $\Sigma_{t}$; injectivity on $(\overline{\Sigma}_{t})^{c}$ still holds when $\gamma=t.$
Continuity of $\Psi_{t,\gamma}=f_{\gamma}$ outside $\overline{\Sigma}_{t}$ follows from [18, Equation (3.8)]. This formula also allows computation of the limiting value as we approach the boundary of $\Sigma_{t}.$ Meanwhile, continuity of $\Psi_{t,\gamma}$ on $\overline{\Sigma}_{t}$ follows from the explicit formulas in [18, Proposition 8.3], which agrees by construction with the limiting value of $\Psi_{t,\gamma}=f_{\gamma}$ on the boundary. ∎
@@
The proof follows the proof of Proposition 4.13 in the additive case. We use the Hamilton–Jacobi analysis for the PDE (5.7) as developed in [18, Section 5], but keeping in mind that $\varepsilon$ in [18] corresponds to $\sqrt{\varepsilon}$ here. Now, [18] uses $\gamma=t$ (i.e., $\tau=0$ in the notation of [18]) as the initial condition. But since we have already established Theorem 5.4 (corresponding to the case $\gamma=0$), it is convenient to use $\gamma=0$ as our initial condition.
We then have the second Hamilton–Jacobi formula from [18, Theorem 5.1] for the PDE (5.7), adapted to the notation used here:
@@
where the curves $\lambda(\gamma),$ $\varepsilon(\gamma),$ and $p_{\varepsilon}(\gamma)$ are given explicitly in Eqs. (5.6)–(5.9) of [18]. Now, if we take $\gamma=0$ as our initial condition (and use the $\varepsilon$” regularization), then the formulas in [18, Equations (5.4) and (5.5)] for the initial momenta become
@@
Now, if we take $\gamma=0$ as our initial condition, the formula for $\lambda(\gamma)$ at $\varepsilon_{0}=0$ in [18, Proposition 5.6] becomes
@@
where the second equality is from Lemma 5.7. Furthermore, at $\varepsilon_{0}=0,$ we have $\varepsilon(\gamma)=0,$ by [18, Proposition 5.6].
@@
where the quantity in the bottom right corner is easily seen to equal 1, using the explicit formula for $\varepsilon(\gamma)$ in [18, Equation (5.7)]. The rest of the argument proceeds as in the proof of Proposition 4.13, using Lemma 5.7 in place of Lemma 4.14. ∎
@@
When $\gamma\neq t,$ the image of $\overline{\Sigma}_{t}$ under $\Psi_{t,\gamma}$ is the set denoted $\overline{\Sigma}_{s,\tau}$ in [18, Definition 2.5], where $s$ and $\tau$ in [18] correspond to $t$ and $t-\gamma$, respectively, here. See also Figure 4 in [18]. When $\gamma=t$ (corresponding to $\tau=0$ in [18]), the image of $\overline{\Sigma}_{t}$ under $\Psi_{t,\gamma}$ is the support of the law of the unitary element $uu_{t},$ where $u_{t}$ is Biane’s free unitary Brownian motion. In the $\gamma=t$ case, the restriction of $\Psi_{t,\gamma}$ to $\overline{\Sigma}_{t}$ is the map written in [23, Corollary 4.30] as
@@ 5.3. The case of $xb_{t,\gamma}$
with a limiting value of $1/\tilde{p}_{2}(\lambda)$ when $\left|\lambda\right|^{2}\tilde{p}_{0}(\lambda)=\tilde{p}_{2}(\lambda).$ This function describes the lifetime of Hamiltonian system associated to the PDE (5.1), with initial condition given by the regularized log potential of the non-negative element $x,$ in the limit as $\varepsilon_{0}\rightarrow 0.$ See [9, Definition 3.6 and Proposition 3.7].
@@
  • Although the function $T$ is not known to be upper semicontinuous, the set $\Sigma_{t}$ is open [9, Proposition 3.17].
@@
Assume 0 is outside $\overline{\Sigma}_{t}.$ If $\mu_{x}(\{0\})>0,$ then 0 is in the Brown support of $xb_{t}$ by [9, Theorem 3.21] and, therefore, 0 is in the spectrum of $xb_{t}.$ On the other hand, if $\mu_{x}(\{0\})=0$ (and 0 is outside $\overline{\Sigma}_{t}$) then by [9, Corollary 3.13], 0 is outside the spectrum of $x.$ Thus, $x$ is invertible. Since, also, $b_{t}$ is invertible [4, p. 265], $xb_{t}$ is invertible so that 0 is outside the spectrum of $xb_{t}.$
@@
where $J_{x}$ is the Herglotz transform of $x$ as in (5.8). Outside $\overline{\Sigma}_{t},$ the function $f_{\gamma}$ plays the role of the map $\Psi_{t,\gamma}$ from the case of a unitary initial condition. (Compare Lemma 5.7.) According to [9, Proposition 4.14], $f_{\gamma}$ is injective on $(\overline{\Sigma}_{t})^{c}$ and $f_{\gamma}(\lambda)$ tends to infinity as $\lambda$ tends to infinity. The closed set $D_{t,\gamma}$ in [9, Proposition 4.12] is then defined by the condition
@@
The analysis of 0 is similar to the proof of Theorem 5.9, using [9, Proposition 4.13] in place of [9, Theorem 3.21]. If $\lambda$ is a nonzero point outside $\overline{\Sigma}_{t}$ and we assume $T(\lambda)>0,$ then Theorem 5.9 tells us that $\lambda$ is outside the spectrum of $xb_{t}.$ Then the proof of Proposition 5.6 applies, using the injectivity of $f_{\gamma}$ obtained in [9, Proposition 4.14], showing that $f_{\gamma}(\lambda)$ is outside the spectrum of $xb_{t,\gamma}.$

### v2

@@ 5.1. The case of $ub_{t}$
According to [9, Theorem 2.7], $S$ satisfies the PDE
@@
For all $t>0,$ the spectrum of $ub_{t}$ is contained in $\overline{\Sigma}_{t}.$ Thus, since [25, Theorem 4.28] shows that $\mathrm{supp}(\mathrm{Br}_{ub_{t}})=\overline{\Sigma}_{t}$ and since the support of the Brown measure of any element is contained in its spectrum, we conclude that
@@
If we evaluate at $\varepsilon_{0}=0,$ the lifetime of the solution of the Hamiltonian system is $T(\lambda),$ which is greater than $t$ by Lemma 5.3. Then a general result about flows (e.g., the fact that the sets $M_{t}$ in [27, Theorem 9.12] are open) tells us that for $(\lambda_{0},\varepsilon_{0})$ in a neighborhood of $(\lambda,0),$ the lifetime of the solution with initial conditions $(\lambda_{0},\varepsilon_{0})$ will remain greater than t.t. [^1] This point is more subtle than it may appear because the formula for the lifetime of the Hamiltonian system in [9, Proposition 5.9] is only valid for $\varepsilon_{0}\geq 0.$ The difficulty is that when $\varepsilon_{0}<0,$ the blowup time of the whole system may be smaller than the blowup time in the formula for $p_{\varepsilon}(t)$. See Remark 5.10 in [9].
@@
from [9, Equation (5.8)], where the notation means that $p_{\varepsilon}$ is computed using the initial conditions $\lambda_{0}$ and $\varepsilon_{0}.$ Then
@@
As we have noted, $T(\lambda)$ has a removable singularity at $\left|\lambda\right|=1.$ Furthermore, $p_{\varepsilon,0}(\lambda)$ can be infinity but cannot be zero. Thus, the only way $T(\lambda)$ can be infinite is when $\lambda=0.$ Meanwhile, according to Proposition 4.8 in [25], the function $T(\lambda)$ is the decreasing limit as $\varepsilon_{0}\rightarrow 0^{+}$ of a certain function $t_{\ast}(\lambda,\varepsilon_{0})$ and this function is continuous in $\lambda$ for each $\varepsilon_{0}.$ The claimed semicontinuity of $T$ then follows by an elementary result [32, Theorem 15.84] about semicontinuous functions. ∎
@@
The claimed result is stated in Theorem 4.10 of [25]. There is, however, a small gap in the proof, concerning the case $\left|\lambda\right|=1,$ which we fill in here. Since $T(0)=\infty,$ we only consider nonzero $\lambda.$ Now, a point $\lambda$ with $T(\lambda)=t>0$ will be outside $\overline{\Sigma}_{t}$ if and only if $T$ has a weak local minimum at $\lambda$ (meaning that all $\lambda^{\prime}$ in a neighborhood of $\lambda$ have $T(\lambda^{\prime})\geq t,$ so that such points are outside $\Sigma_{t}$). According to [25, Lemma 4.15], the function $T(re^{i\theta}),$ with $\theta$ fixed, is strictly increasing for $1<r<\infty,$ and strictly decreasing for $0<r<1.$ Thus, any possible weak local minimum of $T$ would have to be at a point on the unit circle.
We now consider points on the unit circle that are outside $\overline{\Sigma}_{t}$ and thus (Lemma 5.2) outside $\sigma(u)=\mathrm{supp}(\mu_{u}).$ Putting $\left|\lambda\right|=1$ in (5.4), and following the proof of Proposition 3.5 in [33], we may then compute that
@@
Thus, $1/T$ cannot have a weak local maximum at $e^{i\theta}$ and $T$ cannot have a weak local minimum at $e^{i\theta}.$ (Note, however, that $T$ can have a weak local minimum on the unit circle, namely when it is zero—in which case, (5.6) becomes meaningless—but this cannot happen at points in the unit circle outside $\mathrm{supp}(\mu_{u}).$) ∎
@@ 5.2. The case of $ub_{t,\gamma}$
where $b_{t}$ is the value of $b_{t,\gamma}$ at $\gamma=0.$ By the $\tau=s$ case of [20], this map agrees with the one denoted $\Phi_{s,\tau}$ in [20, Section 8]. The following result shows that the “model deformation phenomenon” holds in this setting. That is to say, as we vary $\gamma$ with $u$ and $t$ fixed, the Brown measure of $ub_{t,\gamma}$ changes in a very specific way, namely by push-forward under the map $\Psi_{t,\gamma}.$
@@
This result is due to Ho–Zhong [25, Corollary 4.30] in the case $\gamma=0$ and to Hall–Ho [20, Theorem 8.2] for $\gamma\neq 0.$
@@
Thus, $\Psi_{t,\gamma}$ coincides on $(\overline{\Sigma}_{t})^{c}$ with the holomorphic function denoted $f_{\gamma}$ in [20, Definition 2.2]. Furthermore, $\Psi_{t,\gamma}$ is defined and continuous on $\mathbb{C}.$
@@
The formula (5.10) follows from the $\tau=s$ case of [20, Theorem 6.1]. Once (5.10) is established, we see that $\Psi_{t,\gamma}$ coincides on $(\overline{\Sigma}_{t})^{c}$ with the function denoted $f_{s-\tau}$ in [20], where $s$ and $\tau$ in [20] correspond to $t$ and $t-\gamma$, respectively, here. Then the claimed injectivity follows from Theorem 3.8 in [20]. Note that this theorem assumes $\tau\neq 0,$ which means $\gamma\neq t$ in our current notation, but in light of Lemma 5.3, this assumption is only needed to ensure injectivity of $f_{s-\tau}$ on the boundary of $\Sigma_{t}$; injectivity on $(\overline{\Sigma}_{t})^{c}$ still holds when $\gamma=t.$
Continuity of $\Psi_{t,\gamma}=f_{\gamma}$ outside $\overline{\Sigma}_{t}$ follows from [20, Equation (3.8)]. This formula also allows computation of the limiting value as we approach the boundary of $\Sigma_{t}.$ Meanwhile, continuity of $\Psi_{t,\gamma}$ on $\overline{\Sigma}_{t}$ follows from the explicit formulas in [20, Proposition 8.3], which agrees by construction with the limiting value of $\Psi_{t,\gamma}=f_{\gamma}$ on the boundary. ∎
@@
The proof follows the proof of Proposition 4.13 in the additive case. We use the Hamilton–Jacobi analysis for the PDE (5.7) as developed in [20, Section 5], but keeping in mind that $\varepsilon$ in [20] corresponds to $\sqrt{\varepsilon}$ here. Now, [20] uses $\gamma=t$ (i.e., $\tau=0$ in the notation of [20]) as the initial condition. But since we have already established Theorem 5.4 (corresponding to the case $\gamma=0$), it is convenient to use $\gamma=0$ as our initial condition.
We then have the second Hamilton–Jacobi formula from [20, Theorem 5.1] for the PDE (5.7), adapted to the notation used here:
@@
where the curves $\lambda(\gamma),$ $\varepsilon(\gamma),$ and $p_{\varepsilon}(\gamma)$ are given explicitly in Eqs. (5.6)–(5.9) of [20]. Now, if we take $\gamma=0$ as our initial condition (and use the $\varepsilon$” regularization), then the formulas in [20, Equations (5.4) and (5.5)] for the initial momenta become
@@
Now, if we take $\gamma=0$ as our initial condition, the formula for $\lambda(\gamma)$ at $\varepsilon_{0}=0$ in [20, Proposition 5.6] becomes
@@
where the second equality is from Lemma 5.7. Furthermore, at $\varepsilon_{0}=0,$ we have $\varepsilon(\gamma)=0,$ by [20, Proposition 5.6].
@@
where the quantity in the bottom right corner is easily seen to equal 1, using the explicit formula for $\varepsilon(\gamma)$ in [20, Equation (5.7)]. The rest of the argument proceeds as in the proof of Proposition 4.13, using Lemma 5.7 in place of Lemma 4.14. ∎
@@
When $\gamma\neq t,$ the image of $\overline{\Sigma}_{t}$ under $\Psi_{t,\gamma}$ is the set denoted $\overline{\Sigma}_{s,\tau}$ in [20, Definition 2.5], where $s$ and $\tau$ in [20] correspond to $t$ and $t-\gamma$, respectively, here. See also Figure 4 in [20]. When $\gamma=t$ (corresponding to $\tau=0$ in [20]), the image of $\overline{\Sigma}_{t}$ under $\Psi_{t,\gamma}$ is the support of the law of the unitary element $uu_{t},$ where $u_{t}$ is Biane’s free unitary Brownian motion. In the $\gamma=t$ case, the restriction of $\Psi_{t,\gamma}$ to $\overline{\Sigma}_{t}$ is the map written in [25, Corollary 4.30] as
@@ 5.3. The case of $xb_{t,\gamma}$
with a limiting value of $1/\tilde{p}_{2}(\lambda)$ when $\left|\lambda\right|^{2}\tilde{p}_{0}(\lambda)=\tilde{p}_{2}(\lambda).$ This function describes the lifetime of Hamiltonian system associated to the PDE (5.1), with initial condition given by the regularized log potential of the non-negative element $x,$ in the limit as $\varepsilon_{0}\rightarrow 0.$ See [10, Definition 3.6 and Proposition 3.7].
@@
  • Although the function $T$ is not known to be upper semicontinuous, the set $\Sigma_{t}$ is open [10, Proposition 3.17].
@@
Assume 0 is outside $\overline{\Sigma}_{t}.$ If $\mu_{x}(\{0\})>0,$ then 0 is in the Brown support of $xb_{t}$ by [10, Theorem 3.21] and, therefore, 0 is in the spectrum of $xb_{t}.$ On the other hand, if $\mu_{x}(\{0\})=0$ (and 0 is outside $\overline{\Sigma}_{t}$) then by [10, Corollary 3.13], 0 is outside the spectrum of $x.$ Thus, $x$ is invertible. Since, also, $b_{t}$ is invertible [4, p. 265], $xb_{t}$ is invertible so that 0 is outside the spectrum of $xb_{t}.$
@@
where $J_{x}$ is the Herglotz transform of $x$ as in (5.8). Outside $\overline{\Sigma}_{t},$ the function $f_{\gamma}$ plays the role of the map $\Psi_{t,\gamma}$ from the case of a unitary initial condition. (Compare Lemma 5.7.) According to [10, Proposition 4.14], $f_{\gamma}$ is injective on $(\overline{\Sigma}_{t})^{c}$ and $f_{\gamma}(\lambda)$ tends to infinity as $\lambda$ tends to infinity. The closed set $D_{t,\gamma}$ in [10, Proposition 4.12] is then defined by the condition
@@
The analysis of 0 is similar to the proof of Theorem 5.9, using [10, Proposition 4.13] in place of [10, Theorem 3.21]. If $\lambda$ is a nonzero point outside $\overline{\Sigma}_{t}$ and we assume $T(\lambda)>0,$ then Theorem 5.9 tells us that $\lambda$ is outside the spectrum of $xb_{t}.$ Then the proof of Proposition 5.6 applies, using the injectivity of $f_{\gamma}$ obtained in [10, Proposition 4.14], showing that $f_{\gamma}(\lambda)$ is outside the spectrum of $xb_{t,\gamma}.$